TERAHERTZ SPECTROSCOPY OF CaH (X^{2}Σ^{+}), MgH (X^{2}Σ^{+}), AND ZnH (X^{2}Σ^{+}) : EXTREME HYDRIDE SYNTHESIS

\underline{MATTHEW P. BUCCHINO}, Department of Chemistry and Biochemistry, University of Arizona, Tucson, Arizona 85721; and LUCY M. ZIURYS, Department of Chemistry and Biochemistry, Department of Astronomy, and Steward Observatory, University of Arizona, Tucson, Arizona 85721.

Sub-millimeter spectra of CaH (X^{2}Σ^{+}), MgH (X^{2}Σ^{+}), and ZnH (X^{2}Σ^{+}) have been measured in the sub-mm/THz regime (500−805 GHz) using direct-absorption methods. All species were produced by the reaction of metal vapor with H_{2} under DC discharge in a Broida-type oven. Because of efficient molecule production, spectra of ^{26}MgH, ^{66}ZnH, ^{68}ZnH and 70ZnH could be recorded in their natural abundance as well, with high signal-to-noise. Additionally, the N = 0 → 1 and N = 1 → 2 transitions of ^{25}MgH and ^{67}ZnH were measured. Fermi contact, dipolar, and quadrupole constants for the 25Mg and ^{67}Zn nuclei have been accurately determined for the first time in the gas phase, and are in good agreement with previous ESR studies. The hyperfine constants suggest a substantial amount of covalent character is present in metal hydrides.